• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Subedi

    Log entry time 00:01:36 on September27,2009

    Entry number 291842

    keyword=Shift summary

    Both left and right HRS momentum setting = 3.14159 GeV/c.
    Both left and right HRS angle = 14 degree.
    Beam energy (3 pass) = 3483.1045 MeV.
    Beam current = 98.2 uA.
    Half wave plate: IN
    Slug Number=14

    Shift Crew:
    Evaristo Cisbani (3rd), Abdurahim Rakhman (TO), Ramesh Subedi


    16:00 Smooth production inhereted with an ongoing run 12706.

    16:30 Target IOC rebooted. The reboot was required (after taling to Jian Ping) to possibly eliminate red alarm after some of the beam trips.

    17:43 Compton HV tripped.



    Run Good Number of Pairs
    12706 42.4k HappexNoQweak config
    12707 36.2k HappexNoQweak config
    12708 41.0k HappexNoQweak config
    12709 32.5 HappexNoQweak config, stopped early due to a frozen CODA
    12710 41.9 Happex config
    12711 43.9 Happex config
    12712 43.4 Happex config
    12713 42.0 Happex config